OpenStudy (anonymous):

What is the value of f(x) = 9x when x = -2? a. 1 / 81 b. 81 c. 1 / 18 d. 18

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.

OpenStudy (anonymous):

I have no work cause I do not get it. @goformit100

OpenStudy (anonymous):

-2 * 9 = ?

OpenStudy (the_fizicx99):

Char ^_~

OpenStudy (anonymous):

18?

OpenStudy (anonymous):

Ufufufufu

OpenStudy (anonymous):

so its D?

OpenStudy (anonymous):

Why hello there ^w^

OpenStudy (anonymous):

Yes or no? :3 LOL

OpenStudy (anonymous):

@goformit100 stop spamming that everywhere its annoying and ur not helping

OpenStudy (anonymous):

It would be -18 Hmm...

OpenStudy (anonymous):

Is d really -18?

OpenStudy (anonymous):

There is no negative 18? :C

OpenStudy (anonymous):

Yep >_>

OpenStudy (anonymous):

so its 18? or nah?

OpenStudy (the_fizicx99):

Hmm Are you sure it's not \(\ 9^x \)? Because \(\ \sf 9^{-2} = \dfrac{1}{81}\)

OpenStudy (anonymous):

Nah Depends <_< If 9x was 9^x, then the answer would be 1/81 Or A, In other words

OpenStudy (anonymous):

So, Is It 9x, or 9^x? :)

OpenStudy (anonymous):

MIND READER FE FE

OpenStudy (anonymous):

so its A? XD

OpenStudy (anonymous):

Depends Is 9x really 9^x? Or nah?

OpenStudy (anonymous):

9^2

OpenStudy (anonymous):

Then yes It's A ^w^

OpenStudy (anonymous):



OpenStudy (anonymous):

Because 9^-2 = 1/81 :3

OpenStudy (anonymous):

Yep yep ^w^

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're very welcome ~ ^-^

OpenStudy (the_fizicx99):

o_o You should try to learn yer exponential laws :3

OpenStudy (anonymous):

Ufufufufufu